OpenStudy (anonymous):

Devins T shirt business uses the demand function P= -Q+38 and the supply function P=Q-6. According to these functions what will the equilibrium point be for Devons Tshirt business (in other words, the number of t-shirts sold and the price at which they are sold)

OpenStudy (anonymous):

Will give medals

jimthompson5910 (jim_thompson5910):

equilibrium occurs when the prices are the same and the quantities (demanded and sold) are the same So at equilibrium, P is the same value for both equations which allows us to say... P = Q-6 -Q+38 = Q-6 ... replace P with -Q+38 This replacement is possible since the first equation says P = -Q+38 which tells us that `P` and `-Q+38` are equivalent expressions Now solve for Q and tell me what you get.

OpenStudy (anonymous):

Lol this is probably wrong but I got 32

OpenStudy (anonymous):

It sounds wrong.. I still don't understand it😂😂

jimthompson5910 (jim_thompson5910):

32 is incorrect

jimthompson5910 (jim_thompson5910):

-Q+38 = Q-6 38+6 = Q+Q 44 = 2Q 2Q = 44 Q = ??

OpenStudy (anonymous):

Yes can you explain it differently maybe?

OpenStudy (anonymous):

Umm 22?

OpenStudy (anonymous):

Is the incorrect?

jimthompson5910 (jim_thompson5910):

Q = 22 is correct

OpenStudy (anonymous):

awesome thank you!!

jimthompson5910 (jim_thompson5910):

then you'll use Q = 22 to find P P = Q-6 P = 22-6 and I'm sure you know the rest